# Inference on testing the number of spikes in a high-dimensional generalized spiked Fisher matrix

### v2

@@ 1 Introduction
${\bm{\Sigma}}_{1}=\mathbf{{\bm{\Sigma}}}_{2}+\mathbf{\Delta},$ (1.1)
@@
where $\mathbf{\Delta}$ is a symmetric matrix with low rank. In their study, [23] compared the sample eigenvalues against a threshold in a similar way to [15]. However, when ${\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-1}$ satisfies the condition in (1.1), it inherently has equal non-spiked eigenvalues of 1, thereby restricting its applications to more complex covariance structure. Subsequently, motivated by the hypothesis testing problem $\mathcal{H}_{0}:\mathbf{\Sigma}_{1}=\sigma^{2}\mathbf{\Sigma}_{2}\ \text{v.s.}\ \mathcal{H}_{1}:\mathbf{\Sigma}_{1}=\sigma^{2}\mathbf{\Sigma}_{2}+\mathbf{\Delta}$, [24] explored a more general spiked Fisher matrix, where the non-spiked eigenvalues are still assumed to be identical. They introduced a generic criterion to estimate the number of spikes, later extended by [25] to scenarios where the number of spikes diverges as $p$ goes to infinity. Nevertheless, all of these models are unsuitable for more complex structures characterized by varying non-spiked eigenvalues. This leads to the study of the so-called generalized spiked Fisher matrix, as investigated in [11] and [6], where ${\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-1}$ has the spectral form
$\underbrace{\alpha_{1},\cdots,\alpha_{1}}_{m_{1}},\ldots,\underbrace{\alpha_{K},\cdots,\alpha_{K}}_{m_{K}},\beta_{p,M+1},\cdots,\beta_{p,p}.$ (1.2)
@@
Here, the spectrum in (1.2) is arranged in descending order, where $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$ are spikes with multiplicity $m_{1},m_{2},\cdots,m_{K}$, satisfying $m_{1}+m_{2}+\cdots+m_{K}=M$ with $M$ being a fixed integer. These spikes are significantly larger than the remaining non-spiked eigenvalues $\beta_{p,M+1},\cdots,\beta_{p,p}$.
@@
In this paper, we focus on testing the number of spikes in a generalized spiked Fisher matrix, as described in (1.2), when both the sample size and the dimension of variables proportionally tend to infinity. We first propose a general test statistic that relies on a partial linear spectral statistic, and establish its asymptotic normality under the null hypothesis. This test is then applied to two statistical problems: testing the number of significant variables in a large-dimensional linear regression model and testing the change points in the sequence data. Our approach offers several notable advantages compared to existing works. Firstly, our method exhibits superior generality as it is free of population distributions of samples, and does not rely on the diagonal assumption of the population covariance matrices. Furthermore, the proposed method is more comprehensive with general assumptions of $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{2}$, and encompasses the low-rank case of $\mathbf{\Sigma}_{1}-\mathbf{\Sigma}_{2}$ as a special instance. Finally, beyond theoretical advancements, our practical applications demonstrate commendable performance. In the variable selection problem for large-dimensional linear regression, our test accurately identifies the number of significant variables, exhibiting correct sizes and robust behavior. Additionally, in the change point detection problem, our method consistently achieves higher accuracy across various scenarios compared to existing approaches.
@@ 2.1 Model formulation
${{\bf S}}_{1}=\frac{1}{n_{1}}{\bm{\Sigma}}_{1}^{\frac{1}{2}}{{\bf X}}{{\bf X}}^{*}{{\bm{\Sigma}}}_{1}^{\frac{1}{2}}\quad\text{and}\quad{{\bf S}}_{2}=\frac{1}{n_{2}}{{\bm{\Sigma}}}_{2}^{\frac{1}{2}}{{\bf Y}}{{\bf Y}}^{*}{{\bm{\Sigma}}}_{2}^{\frac{1}{2}}.$ (2.1)
@@
${\mathbf{F}}={\mathbf{R}}_{p}^{*}\widetilde{\mathbf{S}}_{1}{\mathbf{R}}_{p}\widetilde{\mathbf{S}}_{2}^{-1}.$ (2.3)
Here, ${\mathbf{R}}_{p}={\bm{\Sigma}}_{1}^{1/2}{\bm{\Sigma}}_{2}^{-{1/2}}$, $\widetilde{\mathbf{S}}_{1}=n_{1}^{-1}{\mathbf{X}}{\mathbf{X}}^{*}$ and $\widetilde{\mathbf{S}}_{2}={n_{2}}^{-1}{\mathbf{Y}}{\mathbf{Y}}^{*}$ are the standardized sample covariance matrices. Thus, if there is no confusion, we denote the matrix in (2.3) as the Fisher matrix ${\bf F}$ and study it from this point onward.
The eigenvalues of ${\mathbf{R}}_{p}^{*}{\mathbf{R}}_{p}={\bm{\Sigma}}_{2}^{-{1/2}}{\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-{1/2}}$ are equivalent to those in the matrix ${\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-1}$. When the difference between $\bm{\Sigma}_{1}$ and $\bm{\Sigma}_{2}$ is general and does not adhere to the low-rank assumption, the spectrum of ${\mathbf{R}}_{p}^{*}{\mathbf{R}}_{p}$ aligns with the form specified in (1.2) and can be rewritten in descending order as
@@
Let $\beta_{p,j}=\alpha_{k},j\in{\mathcal{J}}_{k}$, where ${\mathcal{J}}_{k}$ denotes the set consisting of the ranks of the $m_{k}$-ple eigenvalue $\alpha_{k}$. Then, $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$ are the population spiked eigenvalues with respective multiplicity $m_{k},k=1,2,\cdots,K$, satisfying $m_{1}+m_{2}+\cdots+m_{K}=M$, where $M$ is a fixed integer. The rest $\beta_{p,M+1},\ldots,\beta_{p,p}$ are non-spiked eigenvalues. Then, we call ${\bf F}$ defined in (2.3) as the so-called generalized spiked Fisher matrix by assuming that these spikes are much larger than the remaining eigenvalues.
@@
For the generalized spiked Fisher matrix in (2.3), let its sample eigenvalues be denoted as $l_{1}({\bf F})\geq l_{2}({\bf F})\geq\cdots\geq l_{p}({\bf F})$. To apply random matrix theory to hypothesis testing problems, we need to establish the central limit theory (CLT) of linear spectral statistics composed of these eigenvalues. To this end, we first focus on their first-order limits. Particularly, for the sample spiked eigenvalues $\{l_{j}({\bf F}),j\in{\mathcal{J}}_{k}\}$, $k=1,2,\ldots,K$ associated with the population spiked eigenvalue $\alpha_{k}$, [11] obtained the following relationship under the separation condition $\min_{i\neq k}|\alpha_{k}/\alpha_{i}-1|>d$ for some constant $d>0$:
$\frac{l_{j}({\bf F})}{\psi_{k}}-1\to 0,a.s.\quad\leavevmode\nobreak\ j\in{\mathcal{J}}_{k},$ (2.4)
@@
$\psi_{k}:=\psi(\alpha_{k})=\frac{\alpha_{k}\left(1-c_{1}\int\displaystyle\frac{t}{t-\alpha_{k}}{\rm d}H(t)\right)}{1+c_{2}\int\displaystyle\frac{\alpha_{k}}{t-\alpha_{k}}{\rm d}H(t)}.$ (2.5)
@@ 2.2 Test on the number of spikes: the general case
$\mathcal{H}_{0}:M=M_{0}\quad\mbox{v.s.}\quad\mathcal{H}_{1}:M\neq M_{0},$ (2.6)
@@
It can be found that $-2\log L$ primarily depends on $\sum_{i=M+1}^{p}l_{i}$ and $\sum_{i=M+1}^{p}\log(l_{i})$. Inspired by this, for the hypothesis (2.6) in the generalized spiked Fisher matrix, we proposed the test statistic
$\sum_{j=1}^{p}f\{l_{j}({\bf F})\}-\sum_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}.$ (2.7)
@@
There has been some related work for the limit behavior of the sample eigenvalues of $\mathbf{F}$. For the linear spectral statistic $\sum_{j=1}^{p}f\{l_{j}({\bf F})\}$, [28] developed its CLT applicable when the general Fisher matrix has no spikes, while [11] focused on the limit of sample spiked eigenvalues $l_{j}({\bf F}),j\in{\mathcal{J}}_{k},k=1,2,\ldots,K$. However, neither [28] nor [11] can address the problem of testing the number of spikes individually. Therefore, we aim to establish the CLT of the partial linear spectral statistic in (2.7) under the null, offering a fresh perspective on this statistical problem.
To develop the asymptotic theory of the statistic in (2.7), we first introduce some intermediate notations. Let $F^{(c_{1},c_{2},H)}$ represent the LSD of the Fisher matrix ${\bf F}$, and $F^{(c_{n1},c_{n2},H_{n})}$ be the counterpart of $F^{(c_{1},c_{2},H)}$ with the substitution of $(c_{1},c_{2},H)$ with $(c_{n1},c_{n2},H_{n})$. Moreover, we denote $m_{c_{1},c_{2}}(z)$ as the Stieltjes transform of $F^{(c_{1},c_{2},H)}$, and $\underline{m}_{c_{1},c_{2}}(z)=-(1-c_{1})z^{-1}+c_{1}m_{c_{1},c_{2}}(z)$. For brevity, the notation $\underline{m}_{c_{1},c_{2}}(z)$ will be simplified to $\underline{m}(z)$. Additionally, let $\underline{m}_{c_{2}}(z)$ satisfy the equation $z=-\{\underline{m}_{c_{2}}(z)\}^{-1}+c_{2}\int\{t+\underline{m}_{c_{2}}(z)\}^{-1}{\rm d}H(t)$. We further define $h^{2}=c_{1}+c_{2}-c_{1}c_{2}$ and $m_{0}(z)=\underline{m}_{c_{2}}\{-\underline{m}(z)\}$. Then, the CLT of the statistic in (2.7) under the null hypothesis is established as follows, with its proof provided in Appendix A.
@@
For the hypothesis testing problem (2.6), suppose that Assumptions 1 to 3 are satisfied. Then, under the null, the test statistic (2.7) follows
@@
$\Rightarrow\mathcal{N}\left(0,1\right),$ (2.8)
@@
The mean and variance terms, $\mu_{f,H}$ and $\nu_{f,H}$, are functions of $m_{0}$ and their expressions can be found in the equations (A.3) and (A.4).
In Theorem 2.1, a notable distinction from the CLT in [28] is the computation of the centering parameter $d_{1}(f,H_{n})$ in (2.9), whose expression becomes more intricate for spiked Fisher matrices. When the function $f(x)$ involved in test statistics and the population ESD $H_{n}(t)$ have simple forms, explicit expressions for the center term can be computed, as demonstrated in the following section. However, when $f(x)$ and $H_{n}(t)$ become complex, the analytical expressions are not readily available. Similar to Remark 4.1 in [28], we propose a numerical procedure in the supplementary material to approximate $d_{1}(f,H_{n})$ for generalized spiked Fisher matrices, thereby facilitating practical implementation and enhancing the applicability of our approach.
@@ 2.3 The low-rank case of $\mathbf{\Sigma}_{1}-\mathbf{\Sigma}_{2}$
Recall the spiked Fisher matrix adhering to the formula (1.1), the difference $\mathbf{\Delta}$ between $\bm{\Sigma}_{1}$ and $\bm{\Sigma}_{2}$ is of low rank. Thus, the spectrum of ${\mathbf{R}}_{p}^{*}{\mathbf{R}}_{p}$ is arranged in descending order as
@@
For the testing problem (2.6), suppose that Assumptions 1 to 3 and the condition $\mathbf{\Sigma}_{1}=\mathbf{\Sigma}_{2}+\mathbf{\Delta}$ are satisfied, then the asymptotic distribution of the test statistic (2.7) can be expressed as
$T_{f,1}$ $=\nu_{f,1}^{-\frac{1}{2}}\left\{\sum\limits_{j=1}^{p}f\{l_{j}({\bf F})\}\!-\!\!\sum\limits_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}\!-\!d_{1}(f,H_{n})\!-\!\mu_{f,1}\right\}\Rightarrow\mathcal{N}\left(0,1\right),$ (2.10)
@@
Example 2.1.
If $f(x)=x$ and $H(t)=\delta_{\{1\}}$, the statistic in (2.8) simplifies to
@@
Example 2.2.
If $f(x)=\log x$ and $H(t)=\delta_{\{1\}}$, the statistic in (2.8) simplifies to
@@
As presented in Corollary 2.1, the mean and variance terms in Examples 2.1 and 2.2 can be computed using the residues. When $f(x)=x$ and $f(x)=\log x$, some results relevant to computing contour integrals can be found in the references [8] and [27], which are utilized in the proofs of the above examples.

### v3

@@ 1 Introduction
${\bm{\Sigma}}_{1}=\mathbf{{\bm{\Sigma}}}_{2}+\mathbf{\Delta},$ (1)
@@
where $\mathbf{\Delta}$ is a symmetric matrix with low rank. In their study, [23] compared the sample eigenvalues against a threshold in a similar way to [15]. However, when ${\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-1}$ satisfies the condition in (1), it inherently has equal non-spiked eigenvalues of 1, thereby restricting its applications to more complex covariance structure. Subsequently, motivated by the hypothesis testing problem $\mathcal{H}_{0}:\mathbf{\Sigma}_{1}=\sigma^{2}\mathbf{\Sigma}_{2}\ \text{v.s.}\ \mathcal{H}_{1}:\mathbf{\Sigma}_{1}=\sigma^{2}\mathbf{\Sigma}_{2}+\mathbf{\Delta}$, [24] explored a more general spiked Fisher matrix, where the non-spiked eigenvalues are still assumed to be identical. They introduced a generic criterion to estimate the number of spikes, later extended by [25] to scenarios where the number of spikes diverges as $p$ goes to infinity. Nevertheless, all of these models are unsuitable for more complex structures characterized by varying non-spiked eigenvalues. This leads to the study of the so-called generalized spiked Fisher matrix, as investigated in [11] and [6], where ${\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-1}$ has the spectral form
$\underbrace{\alpha_{1},\cdots,\alpha_{1}}_{m_{1}},\ldots,\underbrace{\alpha_{K},\cdots,\alpha_{K}}_{m_{K}},\beta_{p,M+1},\cdots,\beta_{p,p}.$ (2)
@@
Here, the spectrum in (2) is arranged in descending order, where $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$ are spikes with multiplicity $m_{1},m_{2},\cdots,m_{K}$, satisfying $m_{1}+m_{2}+\cdots+m_{K}=M$ with $M$ being a fixed integer. These spikes are significantly larger than the remaining non-spiked eigenvalues $\beta_{p,M+1},\cdots,\beta_{p,p}$.
@@
In this paper, we focus on testing the number of spikes in a generalized spiked Fisher matrix, as described in (2), when both the sample size and the dimension of variables proportionally tend to infinity. We first propose a general test statistic that relies on a partial linear spectral statistic, and establish its asymptotic normality under the null hypothesis. This test is then applied to two statistical problems: testing the number of significant variables in a large-dimensional linear regression model and testing the change points in the sequence data. Our approach offers several notable advantages compared to existing works. Firstly, our method exhibits superior generality as it is free of population distributions of samples, and does not rely on the diagonal assumption of the population covariance matrices. Furthermore, the proposed method is more comprehensive with general assumptions of $\mathbf{\Sigma}_{1}$ and $\mathbf{\Sigma}_{2}$, and encompasses the low-rank case of $\mathbf{\Sigma}_{1}-\mathbf{\Sigma}_{2}$ as a special instance. Finally, beyond theoretical advancements, our practical applications demonstrate commendable performance. In the variable selection problem for large-dimensional linear regression, our test accurately identifies the number of significant variables, exhibiting correct sizes and robust behavior. Additionally, in the change point detection problem, our method consistently achieves higher accuracy across various scenarios compared to existing approaches.
@@ 2.1 Model formulation
${{\bf S}}_{1}=\frac{1}{n_{1}}{\bm{\Sigma}}_{1}^{\frac{1}{2}}{{\bf X}}{{\bf X}}^{*}{{\bm{\Sigma}}}_{1}^{\frac{1}{2}}\quad\text{and}\quad{{\bf S}}_{2}=\frac{1}{n_{2}}{{\bm{\Sigma}}}_{2}^{\frac{1}{2}}{{\bf Y}}{{\bf Y}}^{*}{{\bm{\Sigma}}}_{2}^{\frac{1}{2}}.$ (3)
@@
${\mathbf{F}}={\mathbf{R}}_{p}^{*}\widetilde{\mathbf{S}}_{1}{\mathbf{R}}_{p}\widetilde{\mathbf{S}}_{2}^{-1}.$ (5)
Here, ${\mathbf{R}}_{p}={\bm{\Sigma}}_{1}^{1/2}{\bm{\Sigma}}_{2}^{-{1/2}}$, $\widetilde{\mathbf{S}}_{1}=n_{1}^{-1}{\mathbf{X}}{\mathbf{X}}^{*}$ and $\widetilde{\mathbf{S}}_{2}={n_{2}}^{-1}{\mathbf{Y}}{\mathbf{Y}}^{*}$ are the standardized sample covariance matrices. Thus, if there is no confusion, we denote the matrix in (5) as the Fisher matrix ${\bf F}$ and study it from this point onward.
The eigenvalues of ${\mathbf{R}}_{p}^{*}{\mathbf{R}}_{p}={\bm{\Sigma}}_{2}^{-{1/2}}{\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-{1/2}}$ are equivalent to those in the matrix ${\bm{\Sigma}}_{1}{\bm{\Sigma}}_{2}^{-1}$. When the difference between $\bm{\Sigma}_{1}$ and $\bm{\Sigma}_{2}$ is general and does not adhere to the low-rank assumption, the spectrum of ${\mathbf{R}}_{p}^{*}{\mathbf{R}}_{p}$ aligns with the form specified in (2) and can be rewritten in descending order as
@@
Let $\beta_{p,j}=\alpha_{k},j\in{\mathcal{J}}_{k}$, where ${\mathcal{J}}_{k}$ denotes the set consisting of the ranks of the $m_{k}$-ple eigenvalue $\alpha_{k}$. Then, $\alpha_{1},\alpha_{2},\cdots,\alpha_{K}$ are the population spiked eigenvalues with respective multiplicity $m_{k},k=1,2,\cdots,K$, satisfying $m_{1}+m_{2}+\cdots+m_{K}=M$, where $M$ is a fixed integer. The rest $\beta_{p,M+1},\ldots,\beta_{p,p}$ are non-spiked eigenvalues. Then, we call ${\bf F}$ defined in (5) as the so-called generalized spiked Fisher matrix by assuming that these spikes are much larger than the remaining eigenvalues.
@@
For the generalized spiked Fisher matrix in (5), let its sample eigenvalues be denoted as $l_{1}({\bf F})\geq l_{2}({\bf F})\geq\cdots\geq l_{p}({\bf F})$. To apply random matrix theory to hypothesis testing problems, we need to establish the central limit theory (CLT) of linear spectral statistics composed of these eigenvalues. To this end, we first focus on their first-order limits. Particularly, for the sample spiked eigenvalues $\{l_{j}({\bf F}),j\in{\mathcal{J}}_{k}\}$, $k=1,2,\ldots,K$ associated with the population spiked eigenvalue $\alpha_{k}$, [11] obtained the following relationship under the separation condition $\min_{i\neq k}|\alpha_{k}/\alpha_{i}-1|>d$ for some constant $d>0$:
$\frac{l_{j}({\bf F})}{\psi_{k}}-1\to 0,a.s.\quad\leavevmode\nobreak\ j\in{\mathcal{J}}_{k},$ (6)
@@
$\psi_{k}:=\psi(\alpha_{k})=\frac{\alpha_{k}\left(1-c_{1}\int\displaystyle\frac{t}{t-\alpha_{k}}{\rm d}H(t)\right)}{1+c_{2}\int\displaystyle\frac{\alpha_{k}}{t-\alpha_{k}}{\rm d}H(t)}.$ (7)
@@ 2.2 Test on the number of spikes: the general case
$\mathcal{H}_{0}:M=M_{0}\quad\mbox{v.s.}\quad\mathcal{H}_{1}:M\neq M_{0},$ (8)
@@
It can be found that $-2\log L$ primarily depends on $\sum_{i=M+1}^{p}l_{i}$ and $\sum_{i=M+1}^{p}\log(l_{i})$. Inspired by this, for the hypothesis (8) in the generalized spiked Fisher matrix, we proposed the test statistic
$\sum_{j=1}^{p}f\{l_{j}({\bf F})\}-\sum_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}.$ (9)
@@
There has been some related work for the limit behavior of the sample eigenvalues of $\mathbf{F}$. For the linear spectral statistic $\sum_{j=1}^{p}f\{l_{j}({\bf F})\}$, [28] developed its CLT applicable when the general Fisher matrix has no spikes, while [11] focused on the limit of sample spiked eigenvalues $l_{j}({\bf F}),j\in{\mathcal{J}}_{k},k=1,2,\ldots,K$. However, neither [28] nor [11] can address the problem of testing the number of spikes individually. Therefore, we aim to establish the CLT of the partial linear spectral statistic in (9) under the null, offering a fresh perspective on this statistical problem.
To develop the asymptotic theory of the statistic in (9), we first introduce some intermediate notations. Let $F^{(c_{1},c_{2},H)}$ represent the LSD of the Fisher matrix ${\bf F}$, and $F^{(c_{n1},c_{n2},H_{n})}$ be the counterpart of $F^{(c_{1},c_{2},H)}$ with the substitution of $(c_{1},c_{2},H)$ with $(c_{n1},c_{n2},H_{n})$. Moreover, we denote $m_{c_{1},c_{2}}(z)$ as the Stieltjes transform of $F^{(c_{1},c_{2},H)}$, and $\underline{m}_{c_{1},c_{2}}(z)=-(1-c_{1})z^{-1}+c_{1}m_{c_{1},c_{2}}(z)$. For brevity, the notation $\underline{m}_{c_{1},c_{2}}(z)$ will be simplified to $\underline{m}(z)$. Additionally, let $\underline{m}_{c_{2}}(z)$ satisfy the equation $z=-\{\underline{m}_{c_{2}}(z)\}^{-1}+c_{2}\int\{t+\underline{m}_{c_{2}}(z)\}^{-1}{\rm d}H(t)$. We further define $h^{2}=c_{1}+c_{2}-c_{1}c_{2}$ and $m_{0}(z)=\underline{m}_{c_{2}}\{-\underline{m}(z)\}$. Then, the CLT of the statistic in (9) under the null hypothesis is established as follows, with its proof provided in Appendix A.
@@
For the hypothesis testing problem (8), suppose that Assumptions 1 to 3 are satisfied. Then, under the null, the test statistic (9) follows
@@
$\Rightarrow\mathcal{N}\left(0,1\right),$ (10)
@@
The mean and variance terms, $\mu_{f,H}$ and $\nu_{f,H}$, are functions of $m_{0}$ and their expressions can be found in the equations (23) and (24).
In Theorem 2.1, a notable distinction from the CLT in [28] is the computation of the centering parameter $d_{1}(f,H_{n})$ in (11), whose expression becomes more intricate for spiked Fisher matrices. When the function $f(x)$ involved in test statistics and the population ESD $H_{n}(t)$ have simple forms, explicit expressions for the center term can be computed, as demonstrated in the following section. However, when $f(x)$ and $H_{n}(t)$ become complex, the analytical expressions are not readily available. Similar to Remark 4.1 in [28], we propose a numerical procedure in the supplementary material to approximate $d_{1}(f,H_{n})$ for generalized spiked Fisher matrices, thereby facilitating practical implementation and enhancing the applicability of our approach.
@@ 2.3 The low-rank case of $\mathbf{\Sigma}_{1}-\mathbf{\Sigma}_{2}$
Recall the spiked Fisher matrix adhering to the formula (1), the difference $\mathbf{\Delta}$ between $\bm{\Sigma}_{1}$ and $\bm{\Sigma}_{2}$ is of low rank. Thus, the spectrum of ${\mathbf{R}}_{p}^{*}{\mathbf{R}}_{p}$ is arranged in descending order as
@@
For the testing problem (8), suppose that Assumptions 1 to 3 and the condition $\mathbf{\Sigma}_{1}=\mathbf{\Sigma}_{2}+\mathbf{\Delta}$ are satisfied, then the asymptotic distribution of the test statistic (9) can be expressed as
$T_{f,1}$ $=\nu_{f,1}^{-\frac{1}{2}}\left\{\sum\limits_{j=1}^{p}f\{l_{j}({\bf F})\}\!-\!\!\sum\limits_{j\in{\mathcal{J}}_{k},k=1}^{K}f\{l_{j}({\bf F})\}\!-\!d_{1}(f,H_{n})\!-\!\mu_{f,1}\right\}\Rightarrow\mathcal{N}\left(0,1\right),$ (12)
@@
Example 1.
If $f(x)=x$ and $H(t)=\delta_{\{1\}}$, the statistic in (10) simplifies to
@@
Example 2.
If $f(x)=\log x$ and $H(t)=\delta_{\{1\}}$, the statistic in (10) simplifies to
@@
As presented in Corollary 2.1, the mean and variance terms in Examples 1 and 2 can be computed using the residues. When $f(x)=x$ and $f(x)=\log x$, some results relevant to computing contour integrals can be found in the references [8] and [27], which are utilized in the proofs of the above examples.